OpenStudy (anonymous):

Aluminum cylinder: length 1 m, diameter 1 cm Current: 1000 A Conductivity: σ=35 Sm/〖mm〗^2 Linear temperature coefficient: α=3.8∙〖10〗^(-3) 1/K Quadratic temperature coefficient: β=1.3∙〖10〗^(-6) 1/K^2 Heat transfer coefficient to the ambient air: λ=1 K/(Wm^2 ) 1)• Temperature of the aluminum rod with linear temperature dependency? 2).• Temperature of the aluminum rod with quadratic temperature dependency?

OpenStudy (michele_laino):

is the current inside the aluminum rod a cause or is it an effect?

OpenStudy (michele_laino):

if it is a cause, then inside the aluminum rod we have this electric field \({\mathbf{E}}\): \[{\mathbf{j}} = \sigma {\mathbf{E}}\] where \({\mathbf{j}}\) is the current density vector, and \(\sigma\) is the conductivity of aluminum. Furthermore, the power over volume, dissipated in such rod of aluminum is given by the subsequent formula: \[W = \frac{{{{\mathbf{j}}^2}}}{\sigma }\] provided that we are in static condition of electricity

OpenStudy (anonymous):

Hint: The electrical losses of the aluminum rod P_A transferred via heat to the ambiance can be computed with the temperature at the surface ϑ_S, the ambient temperature ϑ_A, and the surface A with the following equation: P_A=(ϑ_S-ϑ_A)/λA

OpenStudy (michele_laino):

so, thermal energy is: \[dQ = Wdt = \frac{{{{\mathbf{j}}^2}}}{\sigma }dt\] where the magnitude of \({\mathbf{j}}\), is: \[\frac{I}{S} = \frac{I}{{\pi {r^2}}}\] and: \(2r=1\;cm\)

OpenStudy (michele_laino):

I meant \(dQ\) is thermal energy over volume, of course